OpenStudy (anonymous):

I have a problem with an elimination problem and I'm not sure what number to chose to multiply the 2 equations with. here is the problem. 2x-14y=30 and -x-6y=24 these are the two equations that I need to use elimination method to solve help plz

OpenStudy (anonymous):

multiply first by 1/2,or change the order of equations and mutiply first by 2

OpenStudy (anonymous):

Its not funny. Ihave a question about a problem.

OpenStudy (ajprincess):

multiply the second equation by 2 and then add the resulting eqn with ur first equation. u will get x eliminated from the eqns.

OpenStudy (anonymous):

2x-14y=30 ( 1) and -x-6y=24 (2) In such questions try to eliminate one of the variable viz. x or y by equalising the coefficients of these. First we multiply equation 1 with (1) and equation 2 with (2) since the coefficient of x is 2 in equation 1 and 1 in equation 2 which give us the following equation 2x-14y=30 and -2x-12y=48 Now adding both equation give us the following : -26y = 78 which gives y = -78/26 = 3 now putting value of y in 1 you get 2x -42 =30 and x = 72/2 = 36